# Electron-ion recombination in composite interactions in liquid xenon

### v1

@@ III Simulation and Results
In this section, we apply the developed recombination model to three composite interaction examples and compare model predictions with experimental measurements. The model can be made fully data-driven, using only experimental data from single-track events as input, but when necessary or convenient we also use xenon yields evaluated by the public version of NEST Szydagis et al. (2025), which is a fit to large ensembles of experimental data and provides xenon response as a continuous function of event energy and electric field.
We first study the double L-shell EC decay of 124Xe, which is a relatively simple example of composite events with two identical tracks. At the LZ WS2024 electric field of 96.5 V/cm, the photon and charge yields for single beta particles at the average xenon L-shell EC energy of 5.2 keV are evaluated to be 41.8 ph/keV and 31.9 e-/keV by the default NEST version. Due to the composite nature of EC decays, the L-shell EC charge yield is suppressed to $\sim$90of that of betas, leading to a measured value of 27.8 e-/keV in LZ Aalbers et al. (2025). With an exciton-to-ion ratio ($N_{ex}/N_{i}$) of 0.06 and a $W$-value of 13.5 eV assumed for ER events Takahashi et al. (1975); Doke et al. (2002); Dahl (2009), the initial number of ions prior to recombination in a L-shell EC track is estimated to be 363, and the average electron-ion recombination probability is calculated using Eq. 2 to be $\bar{r}$=0.60.
@@
The single-L EC recombination probability is then used to simulate the recombination process for the two L-shell tracks in a 124Xe DEC decay. The simulation begins with two ion tracks and two electron clouds; and then each electron is tested for recombination using the re-normalized probability (Eq. 3). It is noteworthy that DEC decays of xenon produce a different nucleus than single EC decays and thus the atomic energy levels are slightly shifted. We follow the calculation in Niţescu et al. (2024) that predicts a double-L EC decay energy of 10.04 keV and adjust the initial ion numbers accordingly, but assume that the track recombination probability remains the same. 10,000 simulations are carried out to produce a distribution of the number of electrons surviving both in-track and cross-track recombination, as shown in Fig. 1 (blue histogram). The predicted mean charge yield of 174$\pm$10 electrons is slightly lower than the LZ best fit of 180$\pm$11 Aalbers et al. (2024a) but the two values agree within uncertainties. The uncertainty on this model prediction only includes that from the measured yield for single-L EC decays, and effects of the assumed $N_{ex}/N_{i}$ value will be discussed later in this section.
@@
The simulation of the two co-located L-shell tracks in a 124Xe DEC decay is similar to the above process except that there are two initially identical ion tracks and twice more electrons. As a result, every time an electron recombines with an ion, the corresponding track’s recombination probability is updated and both track’s probabilities need to be re-normalized. The predicted charge yield distributions for double-L EC decays of 124Xe using the running-$r$ method are also shown in Fig. 1, with the empirical $\bar{r}(N_{i})$ result in cyan and the TIB-based result in red. The mean charge yields are found to be 190$\pm$11 and 195$\pm$11 electrons, respectively, also consistent with the LZ result but slightly higher.
It should be noted that when cross-track recombination is not considered, two L-shell EC decays, with the combined energy scaled to that of double-L EC decays of 124Xe, would produce $\sim$280 electrons (solid vertical line in Fig. 1). The different implementations of our recombination model, using $\sim$700 electrons as the starting point, produce charge yield predictions around 180 electrons near that observed by LZ. The predicted charge yield values are also significantly lower than the NEST-predicted yield of 10.04 keV beta particles (dashed vertical line in Fig. 1), which is suppressed related to the solid vertical line from increased ionization density in the event. The convergence of the different models and the excellent agreement between model predictions and the LZ measurement strongly support the validity of our proposed recombination enhancement model.
@@
All model implementations have an apparent dependence on $N_{ex}/N_{i}$ for single-L EC tracks, which has been assumed to be 0.06 Takahashi et al. (1975); Dahl (2009). However, other estimates of $N_{ex}/N_{i}$ exist in the literature Doke et al. (2002), and NEST further models it as energy dependent with a value between 0 and 0.2 Szydagis et al. (2025). Figure 2 shows the 124Xe DEC yield predictions for $N_{ex}/N_{i}$ values in this range, with the constant-$r$ result shown in blue and the TIB-based running-$r$ result in red. For both methods, varying $N_{ex}/N_{i}$ only changes the model predictions by $\sim$4%, which also remain in good agreement with the LZ best fit. This is due to our approach of fixing the model predictions for constituent tracks at the experimental observed yields. A higher value of $N_{ex}/N_{i}$ reduces the number of initial ions in a track prior to recombination, and thus requires a weaker recombination strength for the track, which, when applied to double-L EC decays, also partially compensates for the smaller initial ion numbers. In addition to reducing the model’s dependence on the assumed $N_{ex}/N_{i}$ value, anchoring single-track predictions to data also mitigates inaccuracies in the chosen $\bar{r}(N_{i})$ functions. This is reflected in the excellent agreement between the running-$r$ results using the empirical $\bar{r}(N_{i})$ function and TIB, as illustrated in Fig. 1. For this reason, only the more convenient TIB function is used for the running-$r$ test here and for the rest of this work.
@@
Lastly, we study the recombination enhancement of co-located NR and ER tracks as in Migdal signals. The ER component in a typical Migdal interaction consists of an atomic vacancy and a Migdal ionization electron. As discussed earlier in this section, the relaxation of an atomic vacancy produces a composite interaction, and a suppressed charge yield of $\sim$90relative to that of beta particles has been measured for L- and M-shell EC decays of xenon Temples et al. (2021); Aalbers et al. (2025). Although EC decays favor inner shell electrons and the Migdal process favors outer shell electrons, the additional Migdal electron makes the average M-shell Migdal ER energy approximately the same as an M-shell EC decay. Therefore, we conveniently approximate M-shell Migdal ER signals as mono-energetic EC events at 1.1 keV with yields measured by LZ Aalbers et al. (2025). We consider NR energies up to 100 keV, which is slightly above the end point of DD neutron scatters with xenon (74 keV). At each NR energy of interest, the in-track and cross-track electron-ion recombination is simulated for monoenergetic NRs and 1.1 keV Migdal ERs to estimate the total charge yield. We note that the $N_{ex}/N_{i}$ value for NRs is different from that of ERs, and has been estimated to be 0.7–1.1 Dahl (2009); Szydagis et al. (2025). For NRs below 5 keV, the charge and light yields evaluated by NEST may lead to zero or even negative recombination probability for some $N_{ex}/N_{i}$ values  [^2] A negative recombination probability is obtained when the initial track ion number estimated using $N_{ex}/N_{i}$ (Eq. 1) is smaller than the observed charge yield. , in which cases we use the NEST charge yield as the initial ion number and assume no recombination. An electric field of 200V/cm is used so the model prediction can be directly compared to data from the Migdal effect search in liquid xenon at LLNL Xu et al. (2024).
@@
Figure 4: Top - The average number of electrons surviving recombination between a 1.1 keV M-shell Migdal ER and monoenergetic NRs, simulated using the constant-$r$ (blue) and running-$r$ (red) methods, in comparison with the simple sum of the signals without considering cross-track recombination, with full (dashed black) and 90% (solid black) beta yield assumed for the Migdal ER component. Bottom - Fractions of the evaluated charge yields relative to the dashed black line in the top figure. An electric field of 200 V/cm is used.
@@
Figure 4 (top) shows the predicted charge yields for ER+NR signals using the constant-$r$ (blue) and TIB-based running-$r$ (red) methods. The two methods produce similar results with a typical difference of $\sim$2%. The predictions are compared to the sum of the NR and ER responses of liquid xenon (ignoring cross-track recombination) with (solid black) and without (dashed black) considering charge quenching in the ER component. Normalized to the ER-NR sum without considering any recombination enhancement, the fractional charge yield slowly decreases with energy to a minimum of $\sim$80between 20 keV and 30 keV. At high NR energies the charge suppression becomes less significant because the ER component and its effect on the total charge output are relatively small. At very low NR energies, the NR component is small relative to the ER which also reduces the effect of cross-track recombination. Varying the $N_{ex}/N_{i}$ value for the NR (ER) component in the full range changes the charge yield estimate by 9% (4%) at 7 keV and by $<$1% (3%) at 100 keV.
@@
Figure 5: Reanalysis of Migdal search data in Xu et al. (2024) using a Migdal signal yield from this work. The electron spectrum with scintillation signals between 4 and 10 photoelectrons is compared with the best-fit signal and background models: Migdal signal at the nominal strength (solid red), single-scatter NR (blue), neutron multi-scatters in xenon (green) and in passive materials (orange), and the sum of all components (grey); the Migdal signal model that simply adds the NR and ER yields without considering recombination enhancement is also shown (dotted red).
@@
We then apply the charge yield prediction to the 7.0$\pm$1.5 keV xenon recoil data used to search for the Migdal effect Xu et al. (2024). Following the approach in Xu et al. (2024), the pure NR and ER yields are estimated using NEST, with the NR model parameters tuned to single-scatter NR data in the Migdal search, and the ER parameters tuned to in-situ gamma calibrations. The Migdal ER model is separately tuned to produce a lower charge yield so the combined ER+NR yield agrees with model predictions in Fig. 4 at 7 keV. Figure 5 shows the best fit of the signal and background models to experimental data, with the Migdal signal strength fixed at the theoretical value. The suppressed charge yield in the updated Migdal model (solid red), relative to that without considering enhanced recombination (dotted red), causes Migdal events to become degenerate with neutron multiscatter backgrounds in the experiment. As a result, the tension between the Migdal prediction and the experimental observation is relieved, confirming that enhanced recombination between co-located NR and ER tracks is a likely explanation of the null Migdal result observed in Xu et al. (2024).
@@
The recombination model developed in this work will play a critical role in future experimental efforts searching for the Migdal effect in liquid xenon. A promising Migdal search approach is to use very low energy NRs. As shown in Fig. 4 (top), the NR+ER recombination enhancement is negligible for NRs below 2 keV, where the black, blue and red curves coincide. As a result, the modeling uncertainty in this regime may also be neglected, so such Migdal searches will bear the least uncertainty from the cross-track recombination effect. In addition, the constant-$r$ method, through its overestimation of recombination enhancement, can guide a Migdal search plan toward a more conservative direction and make a signal claim more convincing. This work also provides a more robust framework for modeling signals in dark matter searches using the Migdal effect.
@@ IV Discussion
The model developed in this work is tested against experimental data containing composite interactions. Although results of the constant-$r$ and running-$r$ methods are interpreted as lower- and upper-bounds on the charge yield of composite events, the predictions are found to converge in these examples and provide strong constraints. In this section, we discuss the subtleties in the model that are not covered in Sec. II.
@@
Figure 6: The simulated reduction of ion number in one track of a double-L EC decay of 124Xe (solid line) and that in a single-L EC decay (dashed line) as electron-ion recombination progresses using the constant-$r$ method. The X axis is the index of electrons tested for recombination, and the initial ion number in the single-L simulation is scaled to half of the double-L case for direct comparison.
@@
The constant-$r$ method can overestimate recombination by ignoring the reduction of recombination probability for an ion track in composite interactions. Cross-track recombination allows more electrons and ions to recombine, and, as the number of ions in the track decreases, its ability to recombine with additional electrons is reduced. Therefore, the overall recombination probability for a track should be lower in a composite event than by itself. However, this effect is relatively small in the examples studied. Figure 6 shows the reduction of ion counts in one of the two tracks in the DEC decay simulation of 124Xe (solid line), in comparison with that in a single-L EC decay with no second track (dashed line). Although both simulations start with the same track recombination probability, the ion number in the double-L EC decay does not decrease as fast as that in the single L EC decay. This is because each electron is competed for by both ion tracks, which reduces the chance for the electron to recombine with either, as reflected by the probability renormalization in Eq. 3. As a result, the final ion counts in the track are not very different in the two scenarios. Because the overall track recombination probability is approximately related to the average ion count throughout the recombination process, the effective $\bar{r}$ value in 124Xe DEC decays would only be modestly reduced from that in single-L EC decays.
@@
The running-$r$ method incorporates the dependence of the recombination probability on the ion count in a track. However, the use of the instant ion count when the electron of interest recombines or escapes neglects the history of the track, and underestimates the recombination probability. In single-track events this effect is compensated for by tuning the model parameter $\mathcal{C}_{t}$ to anchor the model prediction at experimental observation. In composite interactions, however, the further reduction of ion count from cross-track recombination is under-compensated. This effect causes the recombination probability evaluated by the running-$r$ method to be an underestimate and renders the prediction an upper limit on the charge yield of a composite event. Similar to the constant-$r$ method, the bias in the running-$r$ prediction also depends on the additional reduction of ions in composite events (Fig. 6). It is worth noting that a 124Xe DEC decay consists of two identical tracks with high recombination probabilities, so cross-track recombination causes a relatively large change in the final ion count of each track, which leads to a significant difference in the predicted charge yields using the constant-$r$ and running-$r$ methods (Fig. 1). When two tracks with a large disparity in ion numbers co-locate, cross-track recombination may introduce a larger inaccuracy for the smaller track, but its small contribution to the total signal can mitigate the overall effect. In addition, tracks with low ion counts in liquid xenon, such as very low-energy ERs and NRs, usually have low recombination probabilities which helps limit the extra loss of ions to cross-track recombination. For this reason, the model predictions in these cases may provide tighter bounds, as exampled by Fig. 3 and Fig. 4.
@@
In addition to uncertainties from the modeling methods, another significant source of uncertainty is from inaccuracies in the input yields of constituent tracks. For example, a 1 e-/keV shift in the assumed single-L xenon EC yield will change the predicted charge yield of 124Xe DEC decays (10.04 keV) by approximately 10 electrons, comparable to the uncertainty on the LZ best-fit. For the running-$r$ method, an additional uncertainty comes from the function that is used to describe the dependence of the recombination probability on the track ion count. Two functions are studied for 124Xe DEC decays  [^3] The constant-$r$ method may be seen as a special case where the recombination probability does not vary with the track ion count.: one derived empirically using the NEST-evaluated beta yields, and the other using the TIB model. The close match of the two predictions, as shown in Fig. 1, suggests that the uncertainty on either function is relatively small in this energy region, so the analytical TIB function has been used for most of this work.
@@
For recombination in composite events consisting of similar tracks, such as in a 124Xe DEC decay, the TIB model can be used by scaling up the number of total ions in the event from the single-track value. Within the TIB framework, this scaling represents an increase in the ionization density without changing the position distribution. As an approximation, the TIB model can be generalized by adding the values of $\xi$’s in Eq. 4 over all tracks to allow slight variations in the properties of the constituent tracks:
@@
Charge yields estimated using this TIB extrapolation for both 124Xe DEC and Migdal signals are found to be similar to those using the TIB-based running-$r$ method. This TIB extrapolation method is also studied in Ref. Aalbers et al. (2025), which reports similar results. To further study the relationship of the two TIB-based approaches, we compare the predicted charge yields for hypothetical 124Xe DEC-like events where the ion count (Fig. 7, top) or the recombination probability (Fig. 7, bottom) of one constituent track is allowed to vary. When the ion count is varied, in the physical region near the single-L EC energy, the results from different models (red: TIB-based running-$r$, blue: constant-$r$, dashed black: TIB extrapolation in Eq. 6) closely track each other, but at high ion counts Eq. 6 predicts an even lower charge yield than the constant-$r$ method (lower bound in our model). This behavior is consistent with a known limitation of TIB for overestimating recombination above 10 keV beta energy, which is mitigated in Dahl (2009) by incorporating the growth of the track size. In Fig. 7 (bottom), excellent agreement is observed near the physical recombination probability of $\bar{r}$=0.6, but the TIB extrapolation predicts significantly lower charge yield in unphysical regions. Most notably, when one track is assigned $\bar{r}\sim$100%, Eq. 6 predicts $\sim$100recombination for the whole event and causes all electrons and ions to recombine. However, in our model, a hypothetical track with 100% recombination probability ceases to exist after all its ions recombine with all of its electrons, and would have no real effect on the charge yield of other tracks. Similarly a hypothetical track with 0% recombination probability only contributes electrons and its ability to recombine with electrons will not be affected by a co-located track. Therefore, we conclude that, when TIB is used as an input, the running-$r$ model developed in this work is compatible with the TIB extrapolation where TIB is valid, but our model retains prediction power beyond the applicable scope of TIB.

### v2

@@ III Simulation and Results
In this section, we apply the developed recombination model to three composite interaction examples and compare model predictions with experimental measurements. The model can be made fully data-driven, using only experimental data from single-track events as input, but when necessary or convenient we also use xenon yields evaluated by the public version of NEST Szydagis et al. (2025), which is a fit to large ensembles of experimental data and provides xenon response as a continuous function of event energy and electric field. The uncertainty in NEST-predicted ER yield values is estimated to be 10–15%Szydagis et al. (2025).
We first study the double L-shell EC decay of 124Xe, which is a relatively simple example of composite events with two identical tracks. At the LZ WS2024 electric field of 96.5 V/cm, the photon and charge yields for single beta particles at the average xenon L-shell EC energy of 5.2 keV are evaluated to be 41.8 ph/keV and 31.9 e-/keV by the default NEST version. Due to the composite nature of EC decays, the L-shell EC charge yield is suppressed to $\sim$90of that of betas, leading to a measured value of 27.8$\pm$1.0 e-/keV in LZ Aalbers et al. (2025). With an exciton-to-ion ratio ($N_{ex}/N_{i}$) of 0.06 and a $W$-value of 13.5 eV assumed for ER events Takahashi et al. (1975); Doke et al. (2002); Dahl (2009), the initial number of ions prior to recombination in a L-shell EC track is estimated to be 363, and the average electron-ion recombination probability is calculated using Eq. 2 to be $\bar{r}$=0.60.
@@
The single-L EC recombination probability $\bar{r}$ is then used to simulate the recombination process for the two L-shell tracks in a 124Xe DEC decay. The simulation begins with two ion tracks and two electron clouds; and then each electron is tested for recombination using the renormalized probability (Eq. 3). It is noteworthy that DEC decays of xenon produce a different nucleus from single EC decays and thus the atomic energy levels are slightly shifted. We follow the calculation in Niţescu et al. (2024) that predicts a double-L EC decay energy of 10.04 keV and adjust the initial ion numbers accordingly, but assume that the track recombination probability remains the same. 10,000 simulations are carried out to produce a distribution of the number of electrons surviving both in-track and cross-track recombination, as shown in Fig. 2 (blue histogram). The predicted mean charge yield of 174$\pm$10 electrons is slightly lower than the LZ best fit of 180$\pm$11 electrons Aalbers et al. (2024a) but the two values agree within uncertainties. The uncertainty on this model prediction only includes that from the measured yield for single-L EC decays, and effects of the assumed $N_{ex}/N_{i}$ value will be discussed later in this section.
@@
The simulation of the two colocated L-shell tracks in a 124Xe DEC decay is similar to the above process except that there are two initially identical ion tracks and twice more electrons. As a result, every time an electron recombines with an ion, the corresponding track’s recombination probability is updated and both track’s probabilities need to be renormalized. The predicted charge yield distributions for double-L EC decays of 124Xe using the running-$r$ method are also shown in Fig. 2, with the empirical $\bar{r}(N_{i})$ result in cyan and the TIB-based result in red. The mean charge yields are found to be 190$\pm$11 and 195$\pm$11 electrons, respectively, also consistent with the LZ result but slightly higher.
It should be noted that when cross-track recombination is not considered, two L-shell EC decays, with the summed energy scaled to that of double-L EC decays of 124Xe, would produce $\sim$280$\pm$10 electrons (solid vertical line in Fig. 2). The different implementations of our recombination model, using $\sim$700 electrons as the starting point, produce charge yield predictions around 180 electrons near that observed by LZ. The predicted charge yield values are also significantly lower than the NEST-predicted yield of 10.04 keV beta particles (dashed vertical line in Fig. 2), even when a conservative 15% NEST yield uncertainty is considered. The convergence of the different models and the excellent agreement between model predictions and the LZ measurement strongly support the validity of our proposed recombination enhancement model.
@@
All model implementations have an apparent dependence on $N_{ex}/N_{i}$ for single-L EC tracks, which has been assumed to be 0.06 Takahashi et al. (1975); Dahl (2009). However, other estimates of $N_{ex}/N_{i}$ exist in the literature Doke et al. (2002), and NEST further models it as energy dependent with a value between 0 and 0.2 Szydagis et al. (2025). Figure 3 shows the 124Xe DEC yield predictions for $N_{ex}/N_{i}$ values in this range, with the constant-$r$ result shown in blue and the TIB-based running-$r$ result in red. For both methods, varying $N_{ex}/N_{i}$ only changes the model predictions by $\sim$4%, which also remain in good agreement with the LZ best fit. This is due to our approach of fixing the model predictions for constituent tracks at the experimental observed yields. A higher value of $N_{ex}/N_{i}$ reduces the number of initial ions in a track prior to recombination, and thus requires a weaker recombination strength for the track, which, when applied to double-L EC decays, also partially compensates for the smaller initial ion numbers. It is worth noting that reducing the liquid xenon $W$-value has a similar effect to decreasing $N_{ex}/N_{i}$, and choosing $W$=11.5 eV Baudis et al. (2021); Anton et al. (2020) lowers the model predictions by $<$4%. In addition to reducing the model’s dependence on the assumed $N_{ex}/N_{i}$ and $W$ values, anchoring single-track predictions to data also mitigates inaccuracies in the chosen $\bar{r}(N_{i})$ functions. This is reflected in the excellent agreement between the running-$r$ results using the empirical $\bar{r}(N_{i})$ function and TIB, as illustrated in Fig. 2. For this reason, only the more convenient TIB function is used for the running-$r$ test here and for the rest of this work.
@@
Lastly, we study the recombination enhancement of colocated NR and ER tracks as in Migdal signals. The ER component in a typical Migdal interaction consists of an atomic vacancy and a Migdal ionization electron. As discussed earlier in this section, the relaxation of an atomic vacancy produces a composite interaction, and a suppressed charge yield of $\sim$90relative to that of beta particles has been measured for L- and M-shell EC decays of xenon Temples et al. (2021); Aalbers et al. (2025). Although EC decays favor inner shell electrons and the Migdal process favors outer shell electrons, the additional Migdal electron makes the average M-shell Migdal ER energy approximately the same as an M-shell EC decay. Therefore, we conveniently approximate M-shell Migdal ER signals as monoenergetic EC events at 1.1 keV with yields measured by LZ Aalbers et al. (2025). We consider NR energies up to 100 keV, which is slightly above the end point of DD neutron scatters with xenon (74 keV). At each NR energy of interest, the in-track and cross-track electron-ion recombination is simulated for monoenergetic NRs and 1.1 keV Migdal ERs to estimate the total charge yield. We note that the $N_{ex}/N_{i}$ value for NRs is different from that of ERs, and has been estimated to be 0.7–1.1 Dahl (2009); Szydagis et al. (2025). For NRs below 5 keV, the charge and light yields evaluated by NEST may lead to zero or even negative recombination probability for some $N_{ex}/N_{i}$ values  [^2] A negative recombination probability is obtained when the initial track ion number estimated using $N_{ex}/N_{i}$ (Eq. 1) is smaller than the observed charge yield. , in which cases we use the NEST charge yield as the initial ion number and assume no recombination. An electric field of 200V/cm is used so the model prediction can be directly compared to data from the Migdal effect search in liquid xenon at LLNL Xu et al. (2024).
@@
Figure 5: Top: the average number of electrons surviving recombination between a 1.1 keV M-shell Migdal ER and monoenergetic NRs, simulated using the constant-$r$ (blue) and running-$r$ (red) methods, in comparison with the simple sum of the signals without considering cross-track recombination, with full (dashed black) and 90% (solid black) beta yield assumed for the Migdal ER component. Bottom: fractions of the evaluated charge yields relative to the dashed black line in the top figure. An electric field of 200 V/cm is used.
@@
Figure 5 (top) shows the predicted charge yields for NR+ER signals using the constant-$r$ (blue) and TIB-based running-$r$ (red) methods. The two methods produce similar results with a typical difference of $\sim$2%. The predictions are compared to the sum of the NR and ER responses of liquid xenon (ignoring cross-track recombination) with (solid black) and without (dashed black) considering charge suppression in the ER component. Normalized to the NR+ER sum without considering any recombination enhancement, the fractional charge yield slowly decreases with energy to a minimum of $\sim$80between 20 keV and 30 keV. At high NR energies the charge suppression becomes less significant because the ER component and its effect on the total charge output are relatively small. At very low NR energies, the NR component is small relative to the ER which also reduces the effect of cross-track recombination. Varying the $N_{ex}/N_{i}$ value for the NR (ER) component in the full range changes the charge yield estimate by 9% (4%) at 7 keV and by $<$1% (3%) at 100 keV.
@@
Figure 6: Reanalysis of Migdal search data in Xu et al. (2024) using a Migdal signal yield from this work. The electron spectrum with scintillation signals between 4 and 10 photoelectrons is compared with the best-fit signal and background models: Migdal signal at the nominal strength (solid red), single-scatter NR (blue), neutron multiscatters in xenon (green) and in passive materials (orange), and the sum of all components (grey); the Migdal signal model that simply adds the NR and ER yields without considering recombination enhancement is also shown (dotted red).
@@
We then apply the charge yield prediction to the 7.0$\pm$1.5 keV xenon recoil data used to search for the Migdal effect Xu et al. (2024). Following the approach in Xu et al. (2024), the pure NR and ER yields are estimated using NEST, with the NR model parameters tuned to single-scatter NR data in the Migdal search, and the ER parameters tuned to in-situ gamma calibrations. The Migdal ER model is separately tuned to produce a lower charge yield so the combined NR+ER yield agrees with model predictions in Fig. 5 at 7 keV. Figure 6 shows the best fit of the signal and background models to experimental data, with the Migdal signal strength fixed at the theoretical value. The suppressed charge yield in the updated Migdal model (solid red), relative to that without considering enhanced recombination (dotted red), causes Migdal events to become degenerate with neutron multiscatter backgrounds in the experiment. As a result, the tension between the Migdal prediction and the experimental observation is relieved, confirming that enhanced recombination between colocated NR and ER tracks is a likely explanation of the null Migdal result observed in Xu et al. (2024).
@@
The recombination model developed in this work will play a critical role in future experimental efforts searching for the Migdal effect in liquid xenon. A promising Migdal search approach is to use very low energy NRs. As shown in Fig. 5, the NR+ER recombination enhancement is $<$5% below 2 keV, where the modeling uncertainty is also negligible as the blue and red curves coincide. As a result, the modeling uncertainty in this regime may also be neglected, so such Migdal searches will bear the least uncertainty from the cross-track recombination effect. In addition, the constant-$r$ method, through its overestimation of recombination enhancement, can guide a Migdal search plan toward a more conservative direction and make a signal claim more convincing. This work also provides a more robust framework for modeling signals in dark matter searches using the Migdal effect.
@@ IV Discussion
The model developed in this work is tested against experimental data containing composite interactions. Although results of the constant-$r$ and running-$r$ methods are interpreted as lower and upper bounds on the charge yield of composite events, the predictions are found to converge in these examples and provide strong constraints. In this section, we discuss the subtleties in the model that are not covered in Sec. II.
@@
Figure 7: The simulated reduction of ion number in one track of a double-L EC decay of 124Xe (solid line) and that in a single-L EC decay (dashed line) as electron-ion recombination progresses using the constant-$r$ method. The X axis is the index of electrons tested for recombination; the initial ion and electron numbers in the single-L simulation are scaled to half of the double-L case for direct comparison.
@@
The constant-$r$ method can overestimate recombination by ignoring the reduction of recombination probability for an ion track in composite interactions. Cross-track recombination allows more electrons and ions to recombine, and, as the number of ions in the track decreases, its ability to recombine with additional electrons is reduced. Therefore, the overall recombination probability for a track should be lower in a composite event than by itself. However, this effect is relatively small in the examples studied. Figure 7 shows the reduction of ion counts in one of the two tracks in the DEC decay simulation of 124Xe (solid line), in comparison with that in a single-L EC decay with no second track (dashed line). Although both simulations start with the same track recombination probability, the ion number in the double-L EC decay does not decrease as fast as that in the single L EC decay. This is because each electron is competed for by both ion tracks, which reduces the chance for the electron to recombine with either, as reflected by the probability renormalization in Eq. 3. As a result, in both scenarios $\sim$100 ions are left in the track after all electrons available are tested for recombination. Because the overall track recombination probability is approximately related to the average ion count throughout the recombination process, the effective $\bar{r}$ value in 124Xe DEC decays would only be modestly reduced from that in single-L EC decays.
@@
The running-$r$ method incorporates the dependence of the recombination probability on the ion count in a track. However, the use of the instant ion count when the electron of interest recombines or escapes neglects the history of the track, and underestimates the recombination probability. In single-track events this effect is compensated for by tuning the model parameter $\mathcal{C}_{t}$ to anchor the model prediction at experimental observation. In composite interactions, however, the further reduction of ion count from cross-track recombination is undercompensated. This effect causes the recombination probability evaluated by the running-$r$ method to be an underestimate and renders the prediction an upper limit on the charge yield of a composite event. Similar to the constant-$r$ method, the bias in the running-$r$ prediction also depends on the additional reduction of ions in composite events (Fig. 7). It is worth noting that a 124Xe DEC decay consists of two identical tracks with high recombination probabilities, so cross-track recombination causes a relatively large change in the final ion count of each track, which leads to a significant difference in the predicted charge yields using the constant-$r$ and running-$r$ methods (Fig. 2). When two tracks with a large disparity in ion numbers colocate, cross-track recombination may introduce a larger inaccuracy for the smaller track, but its small contribution to the total signal can mitigate the overall effect. In addition, tracks with low ion counts in liquid xenon, such as very low-energy ERs and NRs, usually have low recombination probabilities which helps limit the extra loss of ions to cross-track recombination. For this reason, the model predictions in these cases may provide tighter bounds, as exampled by Fig. 4 and Fig. 5.
@@
In addition to uncertainties from the modeling methods, another significant source of uncertainty is from inaccuracies in the input yields of constituent tracks. For example, a 1 e-/keV shift in the assumed single-L xenon EC yield will change the predicted charge yield of 124Xe DEC decays (10.04 keV) by approximately 10 electrons, comparable to the uncertainty on the LZ best fit. For the running-$r$ method, an additional uncertainty comes from the function that is used to describe the dependence of the recombination probability on the track ion count. Two functions are studied for 124Xe DEC decays  [^3] The constant-$r$ method may be seen as a special case where the recombination probability does not vary with the track ion count.: one derived empirically using the NEST-evaluated beta yields, and the other using the TIB model. The close match of the two predictions, as shown in Fig. 2, suggests that the uncertainty on either function is relatively small in this energy region, so the analytical TIB function has been used for most of this work.
@@
For recombination in composite events consisting of similar tracks, such as in a 124Xe DEC decay, the TIB model can be used by scaling up the number of total ions in the event from the single-track value. Within the TIB framework, this scaling represents an increase in the ionization density without changing the position distribution. As an approximation, the TIB model can be generalized by adding the $\xi$ contribution of each track, $\xi_{k}$, as defined in Eq. 4 to allow slight variations in the properties of the constituent tracks:
@@
Charge yields estimated using this TIB extrapolation for both 124Xe DEC and Migdal signals are found to be similar to those using the TIB-based running-$r$ method. This TIB extrapolation method is also studied in Ref. Aalbers et al. (2025), which reports similar results. To further study the relationship of the two TIB-based approaches, we compare the predicted charge yields for hypothetical 124Xe DEC-like events where the ion count (Fig. 8, top) or the recombination probability (Fig. 8, bottom) of one constituent track is allowed to vary. When the ion count is varied, in the physical region near the single-L EC energy, the results from different models (red: TIB-based running-$r$, blue: constant-$r$, dashed black: TIB extrapolation in Eq. 6) closely track each other. However, at high ion counts Eq. 6 predicts a charge yield that is lower than the constant-$r$ yield prediction, which is interpreted as the lower bound in our model. This behavior is consistent with a known limitation of TIB for overestimating recombination above 10 keV beta energy, which is mitigated in Dahl (2009) by incorporating the growth of the track size. In Fig. 8 (bottom), excellent agreement is observed near the physical recombination probability of $\bar{r}$=0.6, but the TIB extrapolation predicts significantly lower charge yield in unphysical regions. Most notably, when one track is assigned $\bar{r}\sim$100%, Eq. 6 predicts $\sim$100recombination for the whole event and causes all electrons and ions to recombine. However, in our model, a hypothetical track with 100% recombination probability ceases to exist after all its ions recombine with all of its electrons, and would have no real effect on the charge yield of other tracks. Similarly a hypothetical track with 0% recombination probability only contributes electrons and its ability to recombine with electrons will not be affected by a colocated track. Therefore, we conclude that, when TIB is used as an input, the running-$r$ model developed in this work is compatible with the TIB extrapolation where TIB is valid, but our model retains prediction power beyond the applicable scope of TIB.